The sum of my dad's age and my grandfather's age is 163. The difference in their age is 21. Let x= my dad's age and let y= my grandfather's age. Write two equations representing the situation and solve this system using elimination OR substitution.
Answer by stanbon(75887)

The sum of my dad's age and my grandfather's age is 163. The difference in their age is 21. Let x= my dad's age and let y= my grandfather's age. Write two equations representing the situation and solve this system using elimination OR substitution.
------
Equations:
d + g = 163
g - d = 21
------
Add and solve for "g"::
2g = 184
Grandpa's age = 92 yrs
Dad's age = 163-92
----------------
Cheers,
Stan H.
----------